OpenStudy (anonymous):

according to the general equation for conditional probalitiy, if P (A∩B') =1/6 and P(B') =7/18, what is P(A|B)

OpenStudy (anonymous):

\[P(A|B)=\frac{P(A\cap B)}{P(B)}\] you have the numbers you need, plug them in

OpenStudy (anonymous):

oh actually now that i read carefully you don't have those numbers, do you?

OpenStudy (anonymous):

it tells me what they equal so i think i do

OpenStudy (anonymous):

no you have \(P(B')\) but you need \(P(B)\)

OpenStudy (anonymous):

theres a diffence..

OpenStudy (anonymous):

on the other hand, if \(P(B')=\frac{7}{18}\) then you know what \(P(B)\) is right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

does it convert to something

OpenStudy (anonymous):

if something doesn't happen seven out of eighteen times on average, then how often does it happen?

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

right so if \(P(B')=\frac{7}{18}\) then \[P(B)=\frac{11}{18}\]

OpenStudy (anonymous):

another way to say this is that \[P(B')=1-P(B)\]

OpenStudy (anonymous):

the next number you need to know is \(P(A\cap B)\)

OpenStudy (anonymous):

i must be missing something do you have any other information?

OpenStudy (anonymous):

@satellite73 sorry for late reply almost slept

jimthompson5910 (jim_thompson5910):

use the rules of conditional probability to get \[\Large P(A|B') = \frac{P(A \cap B')}{P(B')}\]

OpenStudy (anonymous):

oooh it is \[\Large P(A|B')\]!!!!

OpenStudy (anonymous):

then you do have the numbers you need, ignore all that stuff before and just compute \[\frac{\frac{3}{18}}{\frac{7}{18}}=\frac{3}{7}\]

OpenStudy (anonymous):

oh ok i aww thanks xx

OpenStudy (anonymous):

see*